# High Probability Bound for Cross-Learning Contextual Bandits with Unknown Context Distributions

### v1

@@ B.3 Upper Bounding $\operatorname{\mathbf{bias}}_{4}$
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle(1-F_{e}J_{t})=0$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle\leq 15\sqrt{T\log(\frac{1}{\delta})}$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
@@ B.4.3 Upper Bounding ftrl
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\langle p_{t,c}-\pi_{c},\hat{\ell}_{t,c}\rangle.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\langle p_{t,c}-\pi_{c},\hat{\ell}_{t,c}\rangle$
$\leq$ $\sum_{c}\Pr(c)\left(\frac{1}{\eta}\log K+\frac{\eta}{2}\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\left\langle p_{t,c},\hat{\ell}^{2}_{t,c}\right\rangle\right).$
Here ^$\hat{\ell}_{t,c}^{2}$ denotes the vector formed by squaring each component of ^$\hat{\ell}_{t,c}$.
By Lemma 3, under event $G$, we have ^$\hat{\ell}_{t,c}\leq 2\tilde{\ell}_{t,c}$. Thus assuming event $G$ holds, we can focus on upper bounding
$\sum_{c}\Pr(c)\left(\frac{1}{\eta}\log K+2\eta\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\left\langle p_{t,c},\tilde{\ell}^{2}_{t,c}\right\rangle\right).$
It suffices to upper bound $\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\left\langle p_{t,c},\tilde{\ell}^{2}_{t,c}\right\rangle.$ We have
$\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\left\langle p_{t,c},\tilde{\ell}^{2}_{t,c}\right\rangle$
$=$ $\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}p_{t,c}(a_{t})\tilde{\ell}^{2}_{t,c}(a_{t})$
@@
$\sum_{c}\Pr(c)\left(\frac{1}{\eta}\log K+\frac{\eta}{2}\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\left\langle p_{t,c},\hat{\ell}^{2}_{t,c}\right\rangle\right)$
@@ B.4.4 Upper Bounding $\operatorname{\mathbf{bias}}_{3}$
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\tilde{\ell}_{t,c}\right\rangle.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\tilde{\ell}_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\tilde{\ell}_{t,c}\right\rangle.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
@@
We then bound the second component $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\tilde{\ell}_{t,c}\right\rangle.$ For each time step $t\in\operatorname{\mathcal{T}^{\ell}}$, we define an indicator function
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\tilde{\ell}_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\tilde{\ell}_{t,c}\right\rangle L_{t}.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\tilde{\ell}_{t,c}\right\rangle L_{t}.$
@@
$\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\tilde{\ell}_{t,c}\right\rangle L_{t}$
@@
$\mathbb{E}\left[\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\tilde{\ell}_{t,c}\right\rangle L_{t}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]$
$=$ $\mathbb{E}\left[\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\tilde{\ell}_{t,c}\right\rangle\big{|}\operatorname{\mathcal{H}}_{t-1}\right]L_{t}$
@@
$\left\{\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\tilde{\ell}_{t,c}\right\rangle L_{t}\right\}_{t\in\operatorname{\mathcal{T}^{\ell}}}$
@@
$\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}\right\rangle L_{t}$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\tilde{\ell}_{t,c}\right\rangle L_{t}\leq 4\sqrt{T\log(\frac{1}{\delta})}.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\tilde{\ell}_{t,c}\right\rangle\leq 4\sqrt{T\log(\frac{1}{\delta})}$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\tilde{\ell}_{t,c}\right\rangle\leq 4\sqrt{T\log(\frac{1}{\delta})}+2KT\gamma$
@@ B.5 Combining the pieces
Taking $\iota=2\log(8KT)$, $L=\sqrt{\frac{\iota KT}{\log(K)}}=\widetilde{\Theta}(\sqrt{KT})$, $\gamma=\frac{16\iota}{L}=\widetilde{\Theta}(1/\sqrt{KT})$, and $\eta=\frac{\gamma}{2(2L\gamma+\iota)}=\widetilde{\Theta}(1/\sqrt{KT})$, it is easy to see that $\operatorname{\mathrm{Reg}}(\pi)=\widetilde{\Theta}(\sqrt{KT}\log\frac{1}{\delta})$ with probability at least $1-\frac{1}{KT}-8\delta$ for any policy $\pi$ and any $\delta\in(0,1)$.

### v2

@@ B.3 Upper Bounding $\operatorname{\mathbf{bias}}_{4}$
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle(1-F_{e}J_{t})=0$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle\leq 15\sqrt{T\log(\frac{1}{\delta})}$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
@@ B.4.3 Upper Bounding ftrl
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\langle p_{t,c}-\pi_{c},\widehat{\ell}_{t,c}\rangle.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\langle p_{t,c}-\pi_{c},\widehat{\ell}_{t,c}\rangle$
$\leq$ $\sum_{c}\Pr(c)\left(\frac{1}{\eta}\log K+\frac{\eta}{2}\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\left\langle p_{t,c},\widehat{\ell}^{2}_{t,c}\right\rangle\right).$
Here ^$\widehat{\ell}_{t,c}^{2}$ denotes the vector formed by squaring each component of ^$\widehat{\ell}_{t,c}$.
By Lemma 3, under event $G$, we have ^$\widehat{\ell}_{t,c}\leq 2\widetilde{\ell}_{t,c}$. Thus assuming event $G$ holds, we can focus on upper bounding
$\sum_{c}\Pr(c)\left(\frac{1}{\eta}\log K+2\eta\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\left\langle p_{t,c},\widetilde{\ell}^{2}_{t,c}\right\rangle\right).$
It suffices to upper bound $\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\left\langle p_{t,c},\widetilde{\ell}^{2}_{t,c}\right\rangle.$ We have
$\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\left\langle p_{t,c},\widetilde{\ell}^{2}_{t,c}\right\rangle$
$=$ $\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}p_{t,c}(a_{t})\widetilde{\ell}^{2}_{t,c}(a_{t})$
@@
$\sum_{c}\Pr(c)\left(\frac{1}{\eta}\log K+\frac{\eta}{2}\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\left\langle p_{t,c},\widehat{\ell}^{2}_{t,c}\right\rangle\right)$
@@ B.4.4 Upper Bounding $\operatorname{\mathbf{bias}}_{3}$
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\widetilde{\ell}_{t,c}\right\rangle.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\widetilde{\ell}_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\widetilde{\ell}_{t,c}\right\rangle.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
@@
We then bound the second component $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\widetilde{\ell}_{t,c}\right\rangle.$ For each time step $t\in\operatorname{\mathcal{T}^{\ell}}$, we define an indicator function
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\widetilde{\ell}_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\widetilde{\ell}_{t,c}\right\rangle L_{t}.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\widetilde{\ell}_{t,c}\right\rangle L_{t}.$
@@
$\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\widetilde{\ell}_{t,c}\right\rangle L_{t}$
@@
$\mathbb{E}\left[\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\widetilde{\ell}_{t,c}\right\rangle L_{t}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]$
$=$ $\mathbb{E}\left[\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\widetilde{\ell}_{t,c}\right\rangle\big{|}\operatorname{\mathcal{H}}_{t-1}\right]L_{t}$
@@
$\left\{\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\widetilde{\ell}_{t,c}\right\rangle L_{t}\right\}_{t\in\operatorname{\mathcal{T}^{\ell}}}$
@@
$\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}\right\rangle L_{t}$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\widetilde{\ell}_{t,c}\right\rangle L_{t}\leq 4\sqrt{T\log(\frac{1}{\delta})}.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\widetilde{\ell}_{t,c}\right\rangle\leq 4\sqrt{T\log(\frac{1}{\delta})}$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\widetilde{\ell}_{t,c}\right\rangle\leq 4\sqrt{T\log(\frac{1}{\delta})}+2KT\gamma$
@@ B.5 Combining the pieces
Taking $\iota=2\log(8KT\frac{1}{\delta})$, $L=\sqrt{\frac{\iota KT}{\log(K)}}=\widetilde{\Theta}(\sqrt{KT\log\frac{1}{\delta}})$, $\gamma=\frac{16\iota}{L}=\widetilde{\Theta}(\sqrt{\frac{\log(1/\delta)}{KT}})$, and $\eta=\frac{\gamma}{2(2L\gamma+\iota)}=\widetilde{\Theta}(1/\sqrt{KT\log(1/\delta)})$, it is easy to see that $\operatorname{\mathrm{Reg}}(\pi)=\widetilde{O}(\sqrt{KT\log\frac{1}{\delta}})$ with probability at least $\Pr(G)-8\delta\geq 1-9\delta$ for any policy $\pi$ and any $\delta\in(0,1)$.
The final step is rescaling the probability constant by a factor of $1/9$, which gives that $\operatorname{\mathrm{Reg}}(\pi)=\widetilde{O}(\sqrt{KT\log\frac{1}{\delta}})$ with probability at least $1-\delta$ and ends the proof of Theorem 1.